OpenStudy (anonymous):

What are the solutions of 6x2 - x - 15 = 0?

OpenStudy (anonymous):

do u know the procedure for solving such problems?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok shall i help u

OpenStudy (anonymous):

\[6x^2 - x -15 = 0\] this is the question . just take the coefficient of first and the last term and multiply them

OpenStudy (anonymous):

so the 6 and 15?

OpenStudy (anonymous):

yeah multiply these two terms always remember u have to multiply the co-efficients of first term that contains (variable ^2) and constant term

OpenStudy (anonymous):

on multiplying 6 and 15 what do u get ????????

OpenStudy (anonymous):

90

OpenStudy (anonymous):

factorise it

hero (hero):

-90

OpenStudy (anonymous):

ignore signs

OpenStudy (anonymous):

(x-3) (x+5)

hero (hero):

That's not a good idea

hero (hero):

Just think of two numbers that multiply to get -90 but add to get -1

hero (hero):

-10 and 9

hero (hero):

That's all you gotta do

OpenStudy (anonymous):

no 90 = 2*3*3*5 then u have find such two numbers by mixing them which u will get middle term here we get -10 & 9

OpenStudy (anonymous):

okay now wat

hero (hero):

Sheg what a surprise. You got the same thing I got

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

6x^2 - 10 x + 9 x -15 = 0

OpenStudy (anonymous):

@ hero why surprised dear here in maths u always get the same answers not different

hero (hero):

Now factor by grouping

hero (hero):

6x^2 - 10 x + 9 x -15 = 0 First find the gcf for the first to terms: 2x Then find the gcf for the last two terms: 3 2x(3x-5)-3(3x-5) = 0

hero (hero):

gcf means greatest common factor

hero (hero):

Next use the distributive rule to finish simplifying: ab + ac = a(b+c) In what we are trying to factor, a = 3x-5, b = 2x and c = -3: so 2x(3x-5)-3(3x-5) = 0 becomes (3x-5)(2x-3)

hero (hero):

Make sure you put (3x-5)(2x-3) = 0

hero (hero):

Thanks lil lady

OpenStudy (anonymous):

So whats the answer????